# Extraction of Medication and Temporal Relation from Clinical Text by Harnessing Different Deep Learning Models

### v1

@@ 3.3 Model-II: CNN-BiLSTM for NER
The Convolutional Neural Network (CNN) used in this approach draws inspiration from Chiu and Nichols (2016) hybrid model, which uses bidirectional LSTMs to learn both character and word-level information. As a result, our model employs word embeddings, in addition to other human-created word features and character-level information retrieved using a convolutional neural network. All of these features are sent into a BiLSTM for each word token. Following the tutorial by Hofer (2018), we adapted the code from Chiu and Nichols (2016), to implement the CNN-BiLSTM model for NER task on the 2009 dataset.
@@
The idea of this model is to perform NER prediction by CONCAT of both character-level and word-level features and their combined input knowledge into BiLSTM learning. This enables the model to make better use of character-level features such as prefixes and suffixes, which can reduce the work of manual feature construction. The character-level features are obtained by a CNN model which has been approved as a great success in the NER task and POS tagging task Chotirat and Meesad (2021); Labeau et al. (2015). Per-character feature vectors, such as character embeddings and character kinds, are concatenated with a max layer to produce a new feature vector, which is then used to train a model for recognising words.
@@
The BiLSTM layer forms the core of the network and has the following three inputs: character-level patterns identified by CNN, word-level input from GloVe embeddings, and casing input (word case changes). One of the critical tasks here is feature fusion as mentioned by Chiu and Nichols (2016). Additional features can be applied for example, words can be divided into six categories, including all uppercase letters, all lowercase letters, all numbers, some numbers, and so on. The model includes 7 layers of network structure: character embedding, dropout, 1D convolution, word embedding, concatenation, BiLSTM, and dense out layer (ref. Appendix F: Model Analysis). This CNN-BiLSTM model was set for training for 30 Epoch but convergence after 14 Epoch. Table 5 presents the performance (macro average precision, recall and f1-score) of this model which is better than the scores from the BiLSTM-CRF model. Table 13 shows the parameters used in this model (Appendix).
3.4 Model-III: BERT-CNN for Temporal Relation Extraction
Although the [CLS] tags in BERT output can achieve reasonable classification results, the rich semantic knowledge contained in BERT is not fully utilised. Therefore, in this work, inspired by Michalopoulos et al. (2020), the CNN model is fused to expand BERT learning (structure demonstrated in Figure 24). After the BERT model receives the processed text, the content of the EMRs text is represented by a vector through the two-layer Transformer mechanism of the BERT model. The model produces a vectorised representation of comprehensive semantic information fused by each token vector, sentence vector, and feature vector in the EMRs and then feeds the output of the model to the convolutional neural network. In this work, three different convolutional kernels are used to capture different feature information. After the fully connected layer is connected by word vector mapping, the CNN model further extracts the semantic information of the dialogue text.
3.4.1 Sampling and Evaluation
In order to mitigate the influence of the training dataset imbalance of each sub-class, we down-sampled the dataset to 3000 samples for each class (AFTER, OVERLAP and BEFORE). Table 6 gives the result of the prediction on the training dataset. Table 7 shows the result of $BERT_{base}+CNN$ model performance for the relation extraction (RE) task on the i2b2-2012 test data. This model achieves a macro avg. 64.48% of precision, 67.17% of recall, and 65.03% of F1-score.
Table 6: Temporal RE Evaluation on the 2012 Sampled Training Dataset using BERT-CNN.
category Number / % Number
precision recall f1-score support
AFTER 96.23 97.74 96.86 3000
OVERLAP 93.71 95.85 94.80 3000
BEFORE 97.13 93.24 95.06 3000
accuracy 95.60 9000
macro avg 95.60 95.60 95.60 9000
weighted avg 95.60 95.60 95.60 9000
Table 7: Temporal RE Evaluation on the 2012 Test Dataset using BERT-CNN
category Number / % Number
precision recall f1-score support
AFTER 34.52 49.00 40.52 1122
OVERLAP 67.34 77.23 71.94 4078
BEFORE 91.62 75.28 82.64 6000
accuracy 73.32 11200
macro avg 64.48 67.17 65.03 11200
weighted avg 77.03 73.34 74.47 11200
As the prediction shows, with BERT-base as a word vector extraction or directly as a classification model, the BERT+CNN model can achieve a relatively good classification result, indicating that the pre-trained model BERT-base can extract the semantic information well of dialogue text. Although the BERT+CNN model has reasonable performance for relation extraction, over-fitting on the training dataset is still inevitable, especially for sub-class AFTER and OVERLAP, which show 61.71% and 26.37% lower than the prediction results on the training dataset, respectively. There is still the correlation between the number of test data samples and prediction precision which is that the more supporting samples, the better the prediction performance, which is similar to our NER findings. Finally, our post-processing rules and example outputs are displayed in Appendix (Post-Processing).
4 Related Work
NER and RE belong to the information extraction (IE) task, which has been developed using different methods, from rule-based ones, to machine learning and deep learning models. The rule-based methods require expertise in designing hand-crafted rules suitable for specific tasks and domains. While these methods worked out in some situations, it is often difficult for other researchers to adopt the existing rules into a new experiment Ji et al. (2019). Remarkable work in this category includes Reeves et al. (2012); Strötgen and Gertz (2013); Hao et al. (2018) on temporal expression extractions, and Roberts et al. (2013); Lin et al. (2013) on medical terminology mining.
Earlier stage ML models on TLINKS detection and clinical events and concepts extraction include Chang et al. (2013); Xu et al. (2013); Wu et al. (2015, 2017) where CNNs and RNNs were explored. The combined model of BiLSTM-CRF was also investigated by some existing work reported in Kim and Lee (2020); Lee et al. (2020); Jiang et al. (2019) and in language-specific experiments such as Chinese EMRs by Ji et al. (2019).
Based on these previous findings, our work further investigated various combinations of these state-of-the-art learning models on clinical domain NER and Temporal modelling. More detailed literature can be found in our Appendix (H: More Background and Literatures).
5 Conclusions and Future Work
We carried out empirical study on medication extraction and temporal relation classification using three hybrid machine learning models BiLSTM-CRF, CNN-BiLSTM, and BERT-CNN. By combining these two tasks and a set of rules we designed, we aim to generate a simple structured output indicating if certain medications are in use during the extracted timeframe. While the experimental work demonstrated promising results, we plan to further improve the model performances from different aspects, which include the integration of domain-specific pertained BERT model such as Med-BERT Rasmy et al. (2021), more text features, data-set augmentation e.g. using synthetic data Belkadi et al. (2023), and system optimisation using different learning structures. We also investigate the prompt-based engineering for MedTem2.0 project Cui et al. (2023).
Limitations
Ethical Statement
Appendix
A: Pre-Processing Details
Clinical texts are mixed with different formats. The target of pre-processing is trying to format different data with certain rules and enrich texts with their potential features.
Typically, Pre-processing follows 5 steps: Sentence Segmentation, Tokenisation, Parts of Speech (POS, which explains how a word is used in a sentence), Tagging and Parsing. For our work, GATE Cunningham H (2002), cTAKES Savova et al. (2010) and MedSpaCy (a new clinical text processing toolkit in Python) shows excellent power in building rule-based modules to retrieve extra features. Additionally, the library ’spacy’ is valuable and powerful to link mentions to the associated verb.
Original word Lemmatization POS Tag Explain
The the DET DT det
canning canning NOUN NN compound
process process NOUN NN nsubjpass
for for ADP IN prep
food food NOUN NN pobj
was be AUX VBD auxpass
invented invent VERB VBN ROOT
by by ADP IN agent
Nicolas Nicolas PROPN NNP compound
Appert Appert PROPN NNP pobj
by by ADP IN prep
1999 1999 NUM CD pobj
. . PUNCT . punct
Table 8: POS Tagging Example Marcinkiewicz (1994)
Sentence Boundary Disambiguation
The task of sentence boundary detection is to identify sentences within a text which appear first in an NLP task pipeline, and its performance could decide the following pipeline parts Koehn (2009). Typically, the full stop is detected as the end of a sentence, which is obviously ambiguous and not enough since the full stop is not only used as a sentence boundary delimiter but also juxtaposed with abbreviations (e.g. M.r, A.M, Doc.) which appear quite a lot in clinical texts. Besides, full stop occurs in some special expressions including dates (e.g. 2022.07.22), numbers (e.g. 10.0 mm), and web links (e.g. https://www.google.com/). Sometimes, other delimiters like “!”, “?” could be the end of sentences as well.
In this work, statistical sentence segmentation of Spacy is used and refined with custom rules. For example, given sentence:
The patient ’s bilirubin level at 24 hours of life was 4.6.
Spacy returns three sentences:
  1. 1.
    The patient
  2. 2.
    ’s
  3. 3.
    bilirubin level at 24 hours of life was 4.6
Here we want to make sure that “ ’s ” tokens will never be the start of a sentence. So a rule is added to make sure “ ’s ” never occur at a start of a sentence.
Tokenisation
Having paragraphs segmented into sentences, sentences will be split into tokens, also known as Tokenisation, that is to decompose long texts such as sentences, paragraphs, and articles into word-based data structures for the convenience of subsequent processing and analysis Koehn (2009); Manning and Schutze (1999). Text is some unstructured data; we need first to transform these data into structured data, structured data can be converted into mathematical problems, and Tokenisation is the first step of transformation.
Compared to other script languages (e.g. Chinese), Tokenisation is easier for English corpora since tokens could be separated by space. However, in order to preserve integral information of original texts, endings of contractions ( “’re” in “we’re”, “’t” in “doesn’t”) should receive special treatment (added rules).
POS Tagging
POS Tagging, also called grammatical tagging or word category disambiguation (the POS tags from the Penn Treebank project Marcinkiewicz (1994), which are widely used in practice), refers to the process of marking a word in the corpus as corresponding to a particular part of speech, based on both its definition and context and the output of POS Tagging is a good feature for supervised learning as an input which makes it an important step:
  1. 1.
    POS tagging is important for word sense disambiguation.
  2. 2.
    For example, consider the sentence “They can fish”; One interpretation: They/PRON can/AUX fish/VERB; and another: They/PRON can/VERB fish/NOUN
  3. 3.
    POS tagging is of importance in applications such as machine translation and information retrieval.
Syntactic Parsing
In order to understand sentences, we cannot treat each word in isolation. We need to determine what a word is attached or connected to. Syntactic parsingFrazier (1979) is a technique by which segmented, tokenised, and part-of-speech tagged text is assigned a structure that reveals the relationships between tokens governed by syntax rules, e.g. by grammars. In NLP, rules of context-free grammar (CFG) or probabilistic context-free grammar (PCFG) are used to analyse sentences. Building a parser from scratch is a very complex task in itself. We would pick a grammar like CFG or PCFG, then decide upon which type of parser we want to build. Based on that, we would implement specific algorithms to develop our very own parser. For example, given the sentence “The patient was not started on antibiotics”, the result of Spacy parsing visualisation (Figure 3):
Figure 3: Syntactic Parsing Example using Spacy
B: Word Embeddings
Word embedding techniques play a role in capturing the meaning, semantic relationship, and context of different words while creating word representations Pham and Le (2018). A word embedding is a technique used to generate a dense vector representation of words that include context words about their own. In addition to one-hot embedding that results in sparse matrix embedding, there are enhanced versions of simple bag-of-words models, such as word counts and frequency counters, that essentially represent sparse vectors.
One-hot Encoding
A popular encoding technique is called “one-hot encoding”. In NLP, if a dictionary or sequence has N fields, each field can be represented by an n-dimensional one-hot vector. For example, “The patient was not started on antibiotics” is encoded as in Figure 4 as an example. One-hot encoding can convert classified data into a unified digital format, which is convenient for machine learning algorithms to process and calculate since computers cannot understand non-numeric data. And the fixed dimension vector is convenient for the machine learning algorithm to calculate the linear algebra.
Figure 4: An example of mapping a sentence to a matrix with One-hot Encoding
Data that don’t relate to one another can benefit from one-hot encoding. The arrangement of numbers is treated as a significant characteristic by machine learning algorithms. To put it another way, people will perceive a higher number as superior or more significant than a lower one.
Bag-of-words embedding
A bag-of-words model (BoW) is a way of extracting features from the text for use in modelling, such as with machine learning algorithms, which is a popular way of representing documents Shao et al. (2018). Throughout the realm of information retrieval, the BoW model is used in information retrieval on the assumption that a text may be reduced to a collection of words without regard to its actual order within the document, its grammar, or its syntax. Each word’s presence in the document is autonomous and unrelated to the incidence of other terms. In other words, each word at each place in the document is individually selected without regard to the document’s semantics.
For given two sentences,
  1. 1.
    She is admitted from hospital for a coronary artery bypass surgery.
  2. 2.
    She is a female admitted to outside hospital with chest discomfort.
Based on these two sentences, a dictionary that contains 16 words is built: Dictionary = 1:“She”, 2. “is”, 3. “admitted”, 4. “from”, 5. “hospital”, 6. “for”, 7. “a”, 8. “coronary”, 9. “bypass”, 10. “surgery”, 11. “female”, 12. “to”, 13. “outside”, 14. “with”, 15. “chest”, 16. “discomfort” Then the two sentences can be represented as:
  1. 1.
    $1,1,1,1,1,1,1,1,1,1,1,0,0,0,0,0$
  2. 2.
    $1,1,1,0,1,0,1,0,0,0,1,1,1,1,1,1$
Each element in the vector represents the number of times related words in the dictionary appear in the document. However, in the process of constructing document vectors, we can see that we do not express the order in which words appear in the original sentence.
GloVe Embedding
GloVe stands for “Global Vectors” Pennington et al. (2014). Compared to Word2vec, which only keeps local statistic information, the advantage of GloVe is that it captures both global statistics and local statistics of a corpus, and it is an unsupervised learning pre-trained model that represents words in sentences. It has the ability to extract semantic relationships. The main idea of the Glove method is that “semantic relationships between words can be extracted from the co-occurrence matrix” Bullinaria and Levy (2007). Given a document having a sentence of n words, the co-occurrence matrix X will be a n*n matrix (as in Figure 5) where $X_{ij}$ indicates how many times word i has co-occurred with the word j.
Figure 5: Co-occurrence Matrix with a Window Size of 1
Let $P_{ij}$ represent the likelihood of seeing the words we and j together, which is calculated by dividing the number of times we and j occurred together ($X_{ij}$) by the total number of times word we appeared in the corpus ($X_{i}$), where $P_{ij}=X_{ij}/X_{i}$. Suppose there is a function F that takes in word vectors of we j, and k and returns the desired ratio.
$F(w_{i},w_{j},u_{k})=P_{i}k/P_{j}k$
C: More Details on Learning Models
BiLSTM Illustration
It is very similar to the structure of the traditional recurrent neural networks and multi-layer feed-forward neural networks after expansion, so the problem of gradient disappearance will inevitably occur if there are too many layers during training Hochreiter (1998). The schematic diagram of gradient disappearance is shown in Figure 6. The node colour depth in the figure indicates how much input information at the first moment can be retained at the current moment. The darker the colour is, the more information is retained, and the better the model effect is; the lighter the colour is, the less information is retained, and the model is prone to gradient disappearance. From the feed-forward process of the model, as time goes by, the information that can be extracted at subsequent moments gradually decreases. As shown in the figure (6), when analysing the data from time step 7, it is practically impossible to find the information from time step 1 anymore. When the error of the output layer at time step 7 propagates forward through the gradient during the training process, the model is unable to successfully update the relevant weights at the earlier time because of the gradient’s decline Jiang et al. (2019).
Figure 6: The Vanishing Gradient Problem in Deep Learning Hochreiter (1998)
Figure 7: Standard LSTM Structure Shi et al. (2015)
To illustrate the structure, the BiLSTM model is presented in the Figure 8.
Figure 8: BiLSTM Structure Gan et al. (2021)
Mathematics on CRFs
As an undirected graph model conforming to Markov random field, the conditional distribution $P(Y|X,\lambda)$ of target tag sequence y is calculated based on observation sequence X. Given a clinical text sequence $X=<x_{1},x_{2},...,x_{n}>$, CRFs model aims to make tags on X with a sequence $X=<y_{1},y_{2},...,y_{n}>$ to get this outcome $t_{i}\in\{B,I,E,O\}$, where B E I O represent the start, the end, the inner and the outside of a TIMEx corresponding to $x_{i}$. Given under the condition of random variable X, if the conditional probability distribution of random variable Y $P(Y|X,\lambda)$ that constitutes a conditional random field, which meets the Markov property
$P(Y_{i}|X,Y_{1},Y_{2},...,Y_{n})=P(Y_{i}|X,Y_{i-1},Y_{i+1})$ (1)
The conditional probability can be calculated by the following formula
$P(Y|X)=\frac{1}{Z(X)}e^{\sum_{i=1}^{I}\sum_{k=1}^{K}}\lambda_{k}f_{k}(i,Y_{i},Y_{i-1},X)$ (2)
where Z(X)is the normalisation factor
$Z(X)=\sum_{Y}e^{\sum_{i=1}^{I}\sum_{k=1}^{K}}\lambda_{k}f_{k}(i,Y_{i},Y_{i-1},X)$ (3)
Given the location of the input sequence X and i, where $f_{k}(i,Y_{i},Y_{i-1},X)$ represents the first k eigenvalue values of mark $Y_{i}$ and mark $Y_{i-1}$ where $\lambda_{k}$ represents the feature weights. From the formula, conditional random field model to the algorithm to calculate the sequence through the forward - after the characteristics of different position expectation and conditional probability, then using maximum likelihood estimation of quasi-newton method to find the solution of the model parameters, such as the use of dynamic programming in the Viterbi algorithm for decoding, and the sequence of the output data for testing as shown in the Figure 9. Additionally, there are other possible NLP systems that can be used to extract various features including (1) MedLEE, (2) KnowledgeMap, (3) a Dictionary-based Semantic Tagger (DST).
Figure 9: CRF Structure Lafferty et al. (2001)
Suppose our NER task uses a BiLSTM model with some word embeddings (illustrated in Figure 8). The BiLSTM provides the logits per class for each word, but each prediction for a single word is independent of the predictions for its neighbours. Each token is independently classified without the model knowing the predicted classes of its neighbours.
More on CNNs
Figure 10: Convoluted Neural Networks Topology Saha
To the best of our knowledge, CNN performs admirably on NLP-related tasks like semantic parsing, NER, and Relation Extraction(RE), where deep learning has been efficiently and widely used due to the prevailing interests of academics in this area. ConvNets’ architecture (shown in Figure 10) mimics the human brain’s neuronal connectivity pattern and takes design cues from the Visual Cortex’s structure. Only some areas of the visual field, referred to as the receptive field, are capable of eliciting responses from individual neurons when they are stimulated. A combination of these fields overlaps to encompass the entire visual field. CNN is able to receive an image as input, assign importance (learnable weights and biases) to distinct aspects/objects in the picture, and differentiate between them. Comparatively, ConvNet requires substantially less pre-processing than other classification techniques. In contrast to primitive approaches, ConvNets are capable of learning these filters/characteristics given sufficient training. This concept works just as well in words and characters as it does in practice.
More on BERTs
As discussed in Section H: More Background and Literatures, for downstream tasks modelled by BERT, especially NLP tasks that include sentence text, firstly, the sentences need to be encoded independently and pre-trained through BERT’s self-attention mechanism. It is important to note that in the text classification task, we will need to test the sentence and the label L together and form the text - tag on; this text classification task can be directly into the sentence to the task. Finally, fine-tune the BERT model to find the most suitable parameters for the corresponding task.
Figure 11: BERT Pre-trained Process Shi et al. (2020)
D: Time Expressions and NER Task
Time expressions, medical events, and time relations identifications were performed on 600 clinical notes, and pathological texts from cancer patients Bethard et al. (2016). Conditional Random Fields (CRFs) have been found to be an effective method to detect Time Expressions and events Lin et al. (2013), based on which, we take the CRFs-based method as the recogniser to extract the text spans of TIMEX (e.g. ‘This February’, ‘24th March, 2020’) and EVENT (e.g. ‘Hospitalization from…’, ‘chemotherapy is indicated from… ’). CRFs model is a conditional probability model proposed in 2001 Lafferty et al. (2001). It combines the characteristics of the hidden Markov model (HMM) Eddy (2004) and Maximum Entropy Markov model (MEMM) McCallum et al. (2000), avoids the problem of label bias through global normalisation, and achieves good results in the task of NER Han et al. (2013).
The entity tags used in the i2b2 challenge track 2009 and 2012 have some differences which are shown in Table 9 and Table 10.
Table 9: Tags Used in the i2b2 2009 Dataset.
Tag Meaning Example
m medication Percocet
do dosage 3.2mg
f frequency twice a day
mo mode (/route of administration) Mode: “nm” (not “Tablet”)
du duration 10-day course
r reasons Dizziness
Table 10: Tags Used in the i2b2 2012 Dataset.
Tag Meaning Example
CLINICAL_DEPT clinical department emergency room
EVIDENTIAL events that have an ‘evidential’ nature CT shows
TEST clinical tests CT
PROBLEM symptoms sickness
TREATMENT medications, surgeries and other procedures Levaquin
E: Dataset and Metrics
Despite the popularity and importance of clinical free text mining research, given the sensitivity of patients Schriever (2014), there have not been many public data sets available, and significant challenges are present when researchers are extracting useful information from clinical free text automatically. The EHRs data for this work mostly comes from National NLP Clinical Challenges, “n2c2” formerly called “i2b2”, which are challenges and workshops designed to find ‘the potential of existing clinical records to yield insights that directly impact healthcare improvement’.
To deliver and evaluate the performance of the temporal information extraction system properly, the effectiveness metrics including accuracy, precision, recall, and F-score Han (2014) about the system will be analysed on the training data and test data. To evaluate the machine learning-based classifier, the confusion matrix is widely applied with True Positives indicating an outcome where the model correctly predicts the positive class (TP), True Negatives indicating an outcome where the model correctly predicts the negative class (TN), False Positives indicating an outcome where the model incorrectly predicts the positive class (FP), and False Negatives indicating a result where the model incorrectly predicts the negative category (FN) as measures Manning and Schutze (1999). To make acknowledge whether our system, when finished, would facilitate the work in hospitals, it can be sent to a medical agency to conduct a trial on actual situations and give feedback on it.
$Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$
$Precision=\frac{TP}{TP+FP}$
$Recall=\frac{TP}{TP+FN}$
$F1=\frac{2*Precision*Recall}{Precision+Recall}$
$F1=\frac{2*TP}{2*TP+FP+FN}$
In terms of interpretation, extraction results will be demonstrated in a structured table or CSV shape file with a comprehensive explanation and figure presentation function if necessary. Considering that there are multiple categories of labels, two different kinds of evaluation criteria are applied: 1) macro average evaluation, the function to compute precision, recall and F1-score, for each label and returns the average without considering the proportion for each label in the dataset; and 2) weighted average evaluation, the function to compute precision, recall and F1-score for each label, and returns the average considering the proportion for each label in the dataset.
Data Statistics
The n2c2 dataset (both 2009 and 2012) is composed of clinical notes that have been de-identified. Their annotated formats are not consistent (Figure 12 and Figure 13).
Figure 12: i2b2 2009 Dataset Example
Figure 13: i2b2 2012 Dataset Example
It could be not very pleasant if the chosen model is expecting a different format of training data. A widely used training dataset format is CoNLL Nivre et al. (2007) for NLP NER tasks. In this data presentation, each token (in this case, an individual word or punctuation mark) sits on its own line. The entire tokenised document is presented in that first column. The entity tag is in the last column. A custom parser is required to transform the data from n2c2’s entity-only, offset-based annotation format into CoNLL’s all-token, table-based format. We applied this parser to convert the data from the n2c2 2009 dataset into CoNLL’s format, which data structure consists of sentence row number, separate words (tokens), POS tag and label (as shown in Figure 14). However, there could be slight differences among data from different tracks; this parser could be used on a different dataset.
Figure 14: CoNLL Format
To better understand the dataset, some statistical analysis was taken. As shown in Figure 15, with BIO tagging format, the i2b2 challenge 2009 track dataset labels consist of 13 categories: B-m, I-m, B-do, I-do, B-f, I-f, B-r, I-r, B-mo, I-mo, B-du, and I-du, among which the B-m tag is in the majority of this dataset. A re-sample was conducted to erase the imbalance of the data sample (cutting labels more than 3000 down to 3000).
Figure 15: 2009 Dataset Labels Statistical Analysis
Then, through the statistical analysis of the sentence length of the input text data, the expected value of the normal distribution of the sentence length is taken as the target length of the sentence input in the subsequent model. It can be seen in Figure 16 that the majority of the sentence length is 10-20 words.
Figure 16: 2009 Word Count Statistical Analysis
Similarly, the same work was conducted on the n2c2 challenge 2012 dataset. Most sentences are between 5 and 25 words in length with wider range distribution than the 2009 dataset, as seen in Figure 17.
Figure 17: 2012 Word Count Statistical Analysis
F: Model Analysis
BiLSTM-CRF
Each classification decision is conditionally independent when a simple BiLSTM is followed by a classifier. The linear-chain CRF explicitly describes dependencies between labels as a table, including transition scores between all possible label pairs.
If the labels adhere to a tight internal syntax, the CRF can acquire this information very quickly. There are numerous methods for encoding the output of NER. Still, they typically encode at least the Beginning, Inside, and Outside of an entity, which must be in a syntactically determined order. CRF model will recognise very soon that it is impossible for I-m to follow O, as it must always follow B-m.
As they are conditionally independent, BiLSTM may be uncertain as to whether it should place B-m on a drug or one dosage unit later and, therefore, output both. CRF layer that is aware that this is unlikely and enforces the tags’ intrinsic logic would output B-do, I-do.
LSTM is able to add or remove information to the state unit, which is controlled by structures called gates. Followed by a soft-max layer, although BiLSTM learns adequate context information in the model training stage, the output probabilities of soft-max are independent of each other, and there is no mutual dependence between the output values. Soft-max only selects the label corresponding to the current optimal solution as the output at each step, which cannot avoid the output of wrong labels. For example, arguably, I-m always occurs following B-m instead of B-r. However, with soft-max, this type of apparent mis-classification happens a lot. To solve this problem, following BiLSTM, CRF as the output layer, take account into the order information between labels, so as to obtain more accurate labels through probability calculation, shown in Figure 18.
Figure 18: The Model Structure of BiLSTM+CRF We Re-Implemented
CRF directly models the conditional probability distribution $P(Y|X)$ combining the advantages of graph models to predict multivariate outputs y with a large number of input features x. The linear chain CRF built on Bi-LSTM can effectively model several hard constraints containing the dependencies of output labels.
BiLSTM-CRF Evaluation
The network structure including parameter values of the model for the i2b2 challenge 2009 dataset NER task is shown in Figure 19. Several groups of control experiments are carried out under this model. The main parameters of the final model are shown in Table 11, and the optimal performance on the test set is obtained after 4 epoch training.
Figure 19: Network of BiLSTM Model on the 2009 Dataset
parameter value
DENSE_EMBEDDING 50
LSTM_UNITS 50
LSTM_DROPOUT 0.2
DENSE_UNITS 100
BATCH_SIZE 256
MAX_EPOCHS 10
LEARNING_RATE 0.001
Table 11: Model Parameters of BiLSTM+CRF for the 2009 Dataset
With 70% of the training dataset, 15% validation dataset and 15% test dataset setting, the training process for NER task on the i2b2 2009 dataset is shown in Figure 20 including Loss function and CRF accuracy learning curves.
Figure 20: Training Process on 2009 Dataset
The structure of the model for the i2b2 challenge 2012 dataset NER task is shown in Figure 21. Several groups of control experiments are carried out under this model as well. The main parameters of the final model are shown in Table 12, and the optimal performance on the test set is obtained after 10 epoch training.
Figure 21: Network of BiLSTM Model on the 2012 Dataset
parameter value
DENSE_EMBEDDING 50
LSTM_UNITS 50
LSTM_DROPOUT 0.2
DENSE_UNITS 100
BATCH_SIZE 256
MAX_EPOCHS 30
LEARNING_RATE 0.0001
Table 12: Model Parameters of BiLSTM+CRF for the 2012 Dataset
With 70% of training dataset, 15% validation dataset and 15% test dataset, the training process for NER task on the i2b2 2012 dataset is shown in Figure 22.
Figure 22: Training Process on the 2012 Dataset
CNN-BiLSTM
The seven layers in this model and their functions:
@@ 5 Conclusions and Future Work
The model parameters and structure of CNN-BiLSTM is displayed in Table 13 and Figure 23, with codes adapted from ‘github.com/kamalkraj’ [^2] https://github.com/kamalkraj/Named-Entity-Recognition-with-Bidirectional
-LSTM-CNNs
@@
Table 13: Model Parameters of BiLSTM+CNN for the 2009 Dataset
Figure 23: Network of CNN-BiLSTM Model on the 2009 Dataset Chiu and Nichols (2016)
BERT-CNN
Code adapted from open source projects of UmlsBERT and MedicalRelationExtraction, we trained this BERT+CNN model based on the parameters on the i2b2 challenge 2012 dataset.
The parameters in BERT-CNN model are set as in Table 14 and the model structure is presented in Figure 24.
Figure 24: The Architecture of BERT-CNN Model We Re-Implemented
@@
Table 14: Model Parameters of BERT+CNN
Post-Processing

### v2

@@ III-C Model-II: CNN-BiLSTM for NER
The Convolutional Neural Network (CNN) used in this approach draws inspiration from [32] hybrid model, which uses bidirectional LSTMs to learn both character and word-level information. As a result, our model employs word embeddings, in addition to other human-created word features and character-level information retrieved using a convolutional neural network. All of these features are sent into a BiLSTM for each word token. Following the tutorial by [33], we adapted the code from [32], to implement the CNN-BiLSTM model for NER task on the 2009 dataset.
@@
The idea of this model is to perform NER prediction by CONCAT of both character-level and word-level features and their combined input knowledge into BiLSTM learning. This enables the model to make better use of character-level features such as prefixes and suffixes, which can reduce the work of manual feature construction. The character-level features are obtained by a CNN model which has been approved as a great success in the NER task and POS tagging task [34, 35]. Per-character feature vectors, such as character embeddings and character kinds, are concatenated with a max layer to produce a new feature vector, which is then used to train a model for recognising words.
@@
The BiLSTM layer forms the core of the network and has the following three inputs: character-level patterns identified by CNN, word-level input from GloVe embeddings, and casing input (word case changes). One of the critical tasks here is feature fusion as mentioned by [32]. Additional features can be applied for example, words can be divided into six categories, including all uppercase letters, all lowercase letters, all numbers, some numbers, and so on. The model includes the following 7 layers of network structure and their functions.
@@
The model parameters of CNN-BiLSTM is displayed in Table IX with codes adapted from ‘github.com/kamalkraj’ [^2] https://github.com/kamalkraj/Named-Entity-Recognition-with\-Bidirectional-LSTM-CNNs. This CNN-BiLSTM model was set for training for 30 Epoch but convergence after 14 Epoch. Table VIII presents the performance macro average precision, recall and f1-score of this model which is better than the scores from the BiLSTM-CRF model. The most increase comes from recall score from 52.58 to 77.83 using both Glove embedding, leading to the F1 score increase from 60.86 to 78.17.
TABLE IX: Model Parameters of CNN-BiLSTM for the i2b2 2009 Data
@@
TABLE X: Model Parameters of BERT+CNN
@@
III-D Model-III: BERT-CNN for Temporal Relation Extraction
Figure 5: The Architecture of BERT-CNN Model Re-Implemented
Although the [CLS] tags in BERT output can achieve reasonable classification results, the rich semantic knowledge contained in BERT is not fully utilised. Therefore, in this work, inspired by [36], the CNN model is fused to expand BERT learning. Code adapted from open source projects of UmlsBERT [^3] https://github.com/gmichalo/UmlsBERT and “MedicalRelationExtraction” [^4] https://github.com/chentao1999/MedicalRelationExtraction, we trained this BERT+CNN model based on the parameters on the i2b2 challenge 2012 dataset. The parameters in BERT-CNN model are set as in Table X and the model structure is presented in Figure 5. After the BERT model receives the processed text, the content of the EMRs text is represented by a vector through the two-layer Transformer mechanism of the BERT model. The model produces a vectorised representation of comprehensive semantic information fused by each token vector, sentence vector, and feature vector in the EMRs and then feeds the output of the model to the CNN layer. In this work, three different convolutional kernels are used to capture different feature information. After the fully connected layer is connected by word vector mapping, the CNN model further extracts the semantic information of the dialogue text.
Data Sampling and Evaluation: In order to mitigate the influence of the training dataset imbalance of each sub-class, we down-sampled the dataset to 3000 samples for each class (AFTER, OVERLAP and BEFORE). Table XI gives the result of the prediction on the training dataset. Table XII shows the result of $BERT_{base}+CNN$ model performance for the relation extraction (RE) task on the i2b2-2012 test data. This model achieves a macro avg. 64.48% of precision, 67.17% of recall, and 65.03% of F1-score.
TABLE XI: Temporal RE Evaluation on the 2012 Sampled Training Dataset using BERT-CNN.
category Number / % Number
precision recall f1-score support
AFTER 96.23 97.74 96.86 3000
OVERLAP 93.71 95.85 94.80 3000
BEFORE 97.13 93.24 95.06 3000
accuracy 95.60 9000
macro avg 95.60 95.60 95.60 9000
weighted avg 95.60 95.60 95.60 9000
TABLE XII: Temporal RE Evaluation on the 2012 Test Dataset using BERT-CNN
category Number / % Number
precision recall f1-score support
AFTER 34.52 49.00 40.52 1122
OVERLAP 67.34 77.23 71.94 4078
BEFORE 91.62 75.28 82.64 6000
accuracy 73.32 11200
macro avg 64.48 67.17 65.03 11200
weighted avg 77.03 73.34 74.47 11200
As the prediction shows, with BERT-base as a word vector extraction or directly as a classification model, the BERT+CNN model can achieve a relatively good classification result, indicating that the pre-trained model BERT-base can extract the semantic information well of dialogue text. Although the BERT+CNN model has reasonable performance for relation extraction, over-fitting on the training dataset is still inevitable, especially for sub-class AFTER and OVERLAP, which show 61.71% and 26.37% lower than the prediction results on the training dataset, respectively. There is still the correlation between the number of test data samples and prediction precision which is that the more supporting samples, the better the prediction performance, which is similar to our NER findings.
III-E Post-Processing